OpenStudy (anonymous):

Solve the equation. 1/|2−9x|=1

OpenStudy (alexandervonhumboldt2):

|2-9X|=1

OpenStudy (alexandervonhumboldt2):

2-9X=1 2-9X=-1

OpenStudy (alexandervonhumboldt2):

X=1/9 X=-1/3

OpenStudy (alexandervonhumboldt2):

X HAS TWO SOLUTIONS

Nnesha (nnesha):

@lawls did you understand ???

OpenStudy (anonymous):

yup thanks

Nnesha (nnesha):

okay if you understand then solve this  solve both for x